OpenStudy (anonymous):

can any1 hlp me figure out dis question yet? q is ....a person moves bi car 4rm A to B ,durig journey he covers half of the distance wd speed 60 km/hr find average speed whn 1st 3rd of d rest distance wd speed 72 km /hr and d remainng distance with speed 42 km/hr.

OpenStudy (anonymous):

first of all shruti, pls always type the full question properly as sometimes it is difficult to understand with contraction etc..... Now let the total distance = x km 1/2 of it = x/2 and rest distance = x/2 1/3 of rest distance = 1/3 * x/2 = x/6 final remaining = 2/3 of rest = 2/3*x/6 = 2x/18 = x/9 now time = distance/speed so x/2 time taken for first half x/2, travelling with speed of 60 km/h = T1 = ---- = x/120 h 60 and x/6 time taken for 1/3 of rest, travelling with speed of 72 km/h = T2 = ---- = x/420 h 72 finally x/9 time taken for remaining,travelling with speed of 42 km/h = T3 = ----- = x/378 h 42 so TOTAL TIME OF THE TRIP OF x km = T1 + T2 + T3 x x x 63x + 18x + 20x 101x = --------- + ---------- + ------- h = ----------------- = ------- h 120 420 378 7560 7560 total distance x km 7560 x Average speed = ------------- = -------------- = --------- km/h total time 101x/7560 h 101x = 74.85 km/h

OpenStudy (anonymous):

Harkirat, the process is true, but I think there is a mistake, The last remaining distance sould be \[(2/3)*(x/2)\], so x/3 not x/9, you can verify it by summing up all the distances, consequently, the total time should be \[(281x/15120)\] therefore, the average speed is 53,81 km/h ( I hope I didn't commit any calculation mistake )

OpenStudy (anonymous):

u r correct Rome, i did it yesterday night when i was dead tired, so i made the mistake u pointed out. Thanks n u get a medal for that !!!

OpenStudy (anonymous):

first of all shruti, pls always type the full question properly as sometimes it is difficult to understand with contraction etc..... Now let the total distance = x km 1/2 of it = x/2 and rest distance = x/2 1/3 of rest distance = 1/3 * x/2 = x/6 final remaining = 2/3 of rest = 2/3*x/2 = 2x/6 = x/3 now time = distance/speed so x/2 time taken for first half x/2, travelling with speed of 60 km/h = T1 = ---- = x/120 h 60 and x/6 time taken for 1/3 of rest, travelling with speed of 72 km/h = T2 = ---- = x/420 h 72 finally x/3 time taken for remaining,travelling with speed of 42 km/h = T3 = ----- = x/126 h 42 so TOTAL TIME OF THE TRIP OF x km = T1 + T2 + T3 x x x 21x + 6x + 20x 47x = --------- + ---------- + ------- h = ----------------- = ------- h 120 420 126 2520 2520 total distance x km 2520 x Average speed = ------------- = -------------- = --------- km/h total time 47x/2520 h 47x = 53.62 km/h